### v1

# Positive and Risky Message Assessment for Music Products

Abstract
In this work, we propose a novel research problem: assessing positive and risky messages from music products. We first establish a benchmark for multi-angle multi-level music content assessment and then present an effective multi-task prediction model with ordinality-enforcement to solve this problem. Our result shows the proposed method not only significantly outperforms strong task-specific counterparts but can concurrently evaluate multiple aspects.
1 Introduction
Accessing music has never been more convenient than it is today. People can use various tools, such as high-fidelity players and streaming apps, to enjoy music at any time. Listeners can simply go online, press the PLAY button, and find themselves invigorated after a bad day. However, this easy access also raises concerns that children and adolescents may have a higher chance of being exposed to risky content. Young people’s thoughts and behavior might potentially be affected by the positive or questionable content in songs, as they tend to learn from the modeled behavior that popular music represents Primack et al. (2008). For example, a study has shown an association between adolescent early sexual experiences and degrading sexual content in music (Primack et al., 2009). Similarly, researchers have also revealed that listening to particular types of songs is positively associated with substance use and aggressive behaviors (Chen et al., 2006).
The American Academy of Pediatrics (AAP) holds the opinion that parents should be informed of pediatricians’ concerns regarding potentially harmful effects of music lyrics (American Academy of Pediatrics, 1996). Policymakers have also taken action; for instance, the Recording Industry Association of America (RIAA) introduced the Parental Advisory Label (PAL) program in 1985 to identify audio products with potentially inappropriate content. Such labels were created to draw parents’ attention to products that may not be suitable for their children. Recent studies from the NLP community have made advances in automating the content rating process. Chin et al. and Fell et al. presented machine learning-based methods to automatically classify explicit/non-explicit lyrics in different languages. The learning objectives are derived from content rating systems like PAL, and these works treat the problem as a binary classification task. Despite these existing achievements, however, neither the current PAL system nor the corresponding automated approaches are capable of appraising a music product with diverse information about content suitability.
In this work, we introduce a novel NLP task: assessing the positive and risky messages of a music item. We study the messages that a music item conveys from five significant dimensions regarding appropriateness: Positive Messages, Violence, Substance Consumption, Sex, and Consumerism along three degrees of severity. In the light of those perspectives, we propose a multi-task method that successfully incorporates both content aspect correlations and severity ordinalities to better assess music products. Our research focuses specifically on music performed with vocal techniques (singing, rapping, etc). In such a format of music, the lyrics play a dominant part in conveying the message and opinion from the artists, which has a major impact on the listener’s thoughts and minds. With such an automated approach of music content assessment, any music provider will benefit from service quality improvements for young users and their parents by offering better recommendation or content advisory. Meanwhile, lyricists may also leverage the system to evaluate the portrayal of risky behaviors like ‘get drunk and be somebody’ even at their early stage of creation.
Our contribution: To our best knowledge, this work, for the first time, presents an exploration on assessing music items from both positive and risky aspects on multiple levels. We develop an effective benchmark and made the dataset creation script publicly available for the community. We further propose an effective multi-task ordinality-enforced rating method that can jointly assess different risky aspects with state-of-the-art performance.
2 Music rating and lyrics
To build up a reliable benchmark for assessing positive and risky messages in such lyrics, we exploit expert ratings provided by Common Sense Media (CSM). CSM is a non-profit organization caring for kids’ digital well-being. It hosts a media rating platform that is supported by childhood development experts. The rating system covers various aspects of childhood development and age appropriateness. In this work, based on the focal points from Youth Risk Behavior Surveillance System (YRBSS) Kolbe et al. (1993) introduced by the USA Centers for Disease Control and Prevention, we choose three explicitly risky aspects Violence, Substance Consumption, and Sex. We also pick Consumerism that relates to drawing interest in the acquisition of goods, often associated with a “feel good" experience. Because Consumerism potentially affects the emotional health and identity development in youth Hill (2011). Additionally, we also study Positive Messages in music products in view of the fact that positive values and behaviors are always encouraged in adolescent development.
2.1 Dataset facts and analysis
With CSM expert ratings, we developed a dataset based on publicly available lyrics from the Internet. We collected 1,119 music items (consisting of 10,661 songs) including standard albums, extended plays (EP), long plays (LP), and CD singles. We refer to the collections of songs as albums in this work for simplification. The partition of CD singles and the albums are shown in Table 1. The age recommendation spans from 2 to 18+.
Category Album CD Single
Percentage 62.2% 37.8%
Table 1: Percentage of albums and CD singles.
The duration of individual music items, including both CD singles and albums, shows significant variation. Comprehensive distributions of text lengths are illustrated in Figure 1.
Figure 1: Overall caption for both images
The level of aspect prevalence is from 0 to 5. To reduce class imbalance due to the fact of data availability, we then project such scored ratings into 3-level ordinal categories (Low Presence (0-1), Medium Presence (2-3), High Presence (4-5)) with median split strategy as in related works (Martinez et al., 2019, 2020) from the movie domain. A detailed label distribution is described in Table 2.
Aspect Violence Substance Sex Consumerism Positive
Low 844 743 663 880 736
Medium 190 294 319 209 314
High 85 82 137 30 69
Table 2: Low/medium/high presence item distribution for each message aspect.
We gained permission from CSM to use the music rating data for research, however, by the time we submit this work, the music reviews are no longer shown to the public on CSM’s website until further updates. We still release the names of expert-rated products and links to their lyrics studied in this work for the community.
The music rating data comes from Common Sense Media (CSM). Non-member users can only browse up to three expert product reviews for free each month. We gained permission from CSM to use the music rating data for research, however, by the time we submit this work, the music reviews are no longer shown on CSM’s website until further updates. We still release the names of expert-rated products and links to their lyrics studied in this work for the community. For the lyrics, we do not release them directly for copyright reasons, It is very easy to find lyrics through our provided links and search engines. We make our best effort to collect lyrics from the open Internet but we recognize that there are still very few songs missing.
2.1.1 Inter-dimensional correlation analysis
[…]
Figure 2: Spearman rank correlation between positive and risky prevalence pairs.
We can easily observe from the heat map that typical physical risky behaviors Violence, Substance Consumption, and Sex have a positive correlation with each other, while Consumerism has a positive correlation with those three but less strong. It is no surprise that Positive Messages has a significant negative correlation with three physical risky behaviors. This interdimensional behavior intuitively inspire us to leverage such correlations to design relevant machine learning strategies.
2.1.2 Positive/Risky behavior v.s. Explicitness
In the context of music, explicitness often refers to the content that contains strong language or violence, sex, or substance abuse depictions (from RIAA). It is a generalized description that has a high coincidence with the risky message we studied in this work. Since it is not practical to collect gold labels of explicitness for every music item we studied, we apply an explicit lyrics classifier trained on 438k English lyrics from a previous work Fell et al. (2020). We can further explore the correlation between the explicitness probability and the level of positive and risky messages. Figure 3 illustrates the pattern of explicitness probability towards severity levels of different aspects.
[…]
It is evident that the message ratings are nicely correlated with the explicitness values. For all of the risky message aspects, the explicitness raises as the message ratings go from low to high. Interestingly, the most abstract message, Consumerism, also follows this pattern. On the contrary, Positive Messages also gives a clear negative correlation against explicitness. This correlation study informs us the message ratings for different aspects can be a good complement or alternative to show the inappropriateness for the music items but are more informative, which affirms the significance of our study.
Figure 4: Joint prediction architecture with emotion-guided twin and aspect-aware attention module.
3 Methodology
We formulate this music content assessment problem as a multi-class text classification task, where the ratings of an aspect are the prediction objectives. We take solely lyrics as visible features. Expecting to leverage the correlation pattern among different messages and their ordinal levels, we propose an effective model that incorporates rich semantic representation, aspect-aware multi-task learning, and ordinality-enforcement. The model architecture is illustrated in Figure 4.
Baselines Violence Substance Sex Consumerism Positive Avg
Majority voting 28.65 26.59 24.76 30.22 26.46 27.34
TF-IDF + SVM 51.28 47.31 61.16 40.51 27.79 45.61
BoWV + SVM 46.75 40.23 55.63 31.15 28.28 40.41
TextCNN Kim (2014) 50.59 48.25 59.03 44.88 36.79 47.91
TextRCNN Lai et al. (2015) 57.63 55.86 64.01 44.61 38.47 52.12
LSTM+Attention Shafaei et al. (2020) 33.93 34.48 41.65 34.62 30.06 34.95
BERT Devlin et al. (2019); Fell et al. (2020) 56.93 52.16 60.61 44.96 46.57 52.25
RNN Trans (RT) Zhang et al. (2021) 62.84 62.47 67.36 46.16 44.37 56.64
multi-task + ordinality-enforcement
Soft label 61.42 64.36 68.92 45.21 42.60 56.50
Ranking-classification 65.04 64.41 69.11 45.65 44.72 57.79
Binary transformation 64.46 63.98 69.00 47.36 44.61 57.88
Table 3: Experimental result on positive and risky message level with macro F1 scores with 10-fold cross-validation.
Emotion-guided twin model: We propose to encode the text from two perspectives: universal semantic representation and emotion-centered semantic representation. We employ two pretrained Transformer models finetuned with semantic textual similarity (Sentence-BERT Reimers and Gurevych (2019)) and emotion detection (fine-tuned Distilled RoBERTa on emotion detection tasks Hartmann (2022)) respectively. The final representation is the concatenation of these two types of embeddings.
Joint prediction over multiple aspects: From the dataset study, we found different aspects, specifically risky behaviors, have a potential correlation with each other. For instance, the presence of violence in a music item might appear concurrently with lyrics depicting substance use. The learning objective is a joint loss of predicting ratings of multiple aspects. To fortify the representation uniqueness for different aspects, we further design an aspect-aware attention module to learn specialized features for each aspect. The module begins with a projection layer with non-linearity and further learns weights through an aspect-differentiation matrix. The final representation $\boldsymbol{x}_{out}$ comes from the addition of input $\boldsymbol{x}_{in}$ with residual over skip connections as described in Equation 1.
$\boldsymbol{x}_{out}=\boldsymbol{x}_{in}+\boldsymbol{x}_{in}\circ{Softmax}\left(\boldsymbol{x}_{in}{\boldsymbol{W}}_{attn}\right)$ (1)
Ordinality-enforcement techniques: The severity ratings are discrete interclass ordinal variables instead of independent categorical labels. Typical classification models usually ignore such correlations between ordinal categories. In this work we apply three ordinality-enforcement techniques that are better suited for the task:
Siamese ranking-classification Zhang et al. (2021): This method leverages a Siamese network to process a pair of instances for both ranking and classification objectives. The ranking (comparison) operation has the potential to learn pairwise ordinal differences in severity levels between samples. The model is optimized with two cross-entropy losses for the two objectives as in Equation 2, where $l_{\text{cls}}$ comes from multi-class classification while $l_{\text{rank}}$ is derived from comparing ratings (lower/same/higher) between the two music items.
[…]
Binary attributes transformation Frank and Hall (2001): This method first divide the sorted ordinal label set (with $n$ elements) into two sets by splitting at every two adjacent elements. It is obvious that there are $n-1$ ways of division. Then we further treat each division case as a binary classification problem and predict the ordinal value on the corresponding set. This method applies more than one classifier to exploit the ordering information several times. In the setting of this problem, we split the 3-class ordinal levels for each aspect into two cases considering possible probability intervals:
$\operatorname{Pr}(y_{i}\leqslant\text{Low}),\operatorname{Pr}(y_{i}>\text{Low})\;$ (3)
$\operatorname{Pr}(y_{i}\leqslant\text{Mid}),\operatorname{Pr}(y_{i}>\text{Mid}).$
Therefore, we apply two binary classifiers to those two scenarios.
Soft label Diaz and Marathe (2019): This method introduces a label softening function to convert ordinal category values as discrete distribution. A class label $y_{i}$ from label set is encoded using the following formula into a soft label $c_{i}^{soft}$.
$c_{i}^{soft}=\frac{e^{-\phi\left(y_{t},y_{i}\right)}}{\sum_{k=1}^{K}e^{-\phi\left(y_{t},y_{k}\right)}}\quad\forall y_{i}\in\mathbb{Y}.$ (4)
The Kullback-Leibler divergence is used as loss function to measure the difference between the predicted probability and the soft label. $\phi\left(y_{t},y_{i}\right)=\left|y_{t}-y_{i}\right|$ is chosen as the metric penalty for the sake of simplicity.
All of the selected ordinality-enforcement techniques are applied respectively to the base model with a multi-tasking module.
4 Experiments and Results
Removed sentence from a segment of Heartless (2019) Violence Substance Sex
Rating Low High Mid
Whole segment confidence 98.77 90.57 72.60
1 Stix drunk, but he never miss a target 1.13 -90.56 $\Downarrow$ -64.57 $\downarrow$
2 Photoshoots, I’m a star now (Star) -4.18 2.48 -40.27 $\uparrow$
3 I’m talkin’ Time, Rolling Stone, and Bazaar now (Bazaar now) -1.91 1.55 -15.39
4 Sellin’ dreams to these girls with their guard down (What?) -1.71 2.01 -30.90 $\downarrow$
Table 4: An input perturbation study on the behavior of the proposed ranking-classification model. We choose three risky aspects - Violence, Substance consumption, and Sex - as it performs the best. The numbers indicate the absolute probability change of the original prediction result. A double down arrow $\Downarrow$ indicates the predicted severity downgraded by two levels, a single down arrow $\downarrow$ means downgraded by one, and an up arrow $\uparrow$ represents upgraded by one.
[…]
5 Discussion and analysis
Speaking of aspects, three types of explicit content, Violence, Substance, and Sex are relatively easier for the model to predict than Consumerism and Positive messages. The reason could be that the latter aspects are more abstract concepts comparing to risky behaviors that can be described in the lyrics with an overt manner. For Consumerism, we hypothesize a model that can better capture signals on ’goods’ and encouragement for acquisition may work more effectively. Unlike explicit questionable content like Violence, Positive messages is even harder to intuitively find a clear textual pattern since positive is more heterogeneous. We may need the model to gain a high-level understanding of the content to distinguish the quantity of positive value that a song or an album wants to deliver.
5.1 Ablation study
Ablation Perf change
Best performing model 0.00
Aspect-aware module -0.30
Emotion-guided twin model -0.11
Ordinality enforcement (binary) -0.24
Multitask joint prediction 1.17
Backbone only + multitask -2.52
Table 5: Ablation study of different components in the best-performing model. We report and analyze the average performance changes across the five aspects.
We conducted an ablation study on the best-performing model by iteratively removing individual components from the model architecture. The experiment demonstrated that all modules within the network structure contributed to the performance in prediction. When we configured the proposed network for single-task prediction with all modules, the performance reached even higher levels. This indicates that the effectiveness of the novel network modules can help single-task baseline models learn better quality text representations. Additionally, we experimented with the backbone model by having it predict all five aspects directly in a multitask setting, which resulted in a large performance drop. This finding highlights the effectiveness and necessity of incorporating components that can guide and enhance multitask learning.
5.2 Saliency analysis
We perform saliency analysis using input perturbation to better understand the model prediction behavior. We chose a segment from the lyrics of Heartless (2019) by The Weekend, which contains explicit language. We do the perturbation sentence by sentence through removing one and feeding the rest of the lyrics into the model, then we inspect the model prediction result. Table 4 shows the detailed influence on prediction confidence and outcome. For Substance Consumption, line 1 explicitly contains the word drunk. When we remove this sentence, the prediction drastically changes from High to Low. For Sex, when we remove the first line, the prediction probability decreases and the result becomes Low. The same situation happens to the fourth line. Interestingly, removing line 2 results in an upgrade. We hypothesize that this deletion increases the density of sexual implication content. This case study intuitively shows the model can successfully capture particular mentions of risky messages such as substance use and sex-related topics that have significance in severity prediction.
5.3 Error analysis
Figure 5 shows the confusion matrix of the prediction results of the best-performing model on one-fold. In general, the proposed method can capture the ordinal information well because wrong predictions that cross two levels (predict low to high or high to low) are rare. Specifically, the model struggled in giving correct predictions on high Consumerism. We suppose the number of training instances of high Consumerism is relatively small. The same case happened to high Positive predictions. We hypothesize the heterogeneous nature of Positive content makes it challenging to predict.
Figure 5: Prediction confusion matrix of the best-performing method on all 5 aspects.
5.4 Case study of unsuccessful predictions
We dig into some unsuccessful predictions to analyze the errors. We mainly focus on the hard aspects of the model.
  • Album: The Best Damn Thing (2007) by Avril Lavigne: The proposed model gives Consumerism low and Positive low ratings, however, the correct labels for those two aspects are medium. This album is a pop-punk production and the songs in the album seemed to be targeted at young people with themes such as love and encouragement. For Consumerism, there are explicit lyrics saying:
    I hate it when a guy doesn’t get the tab
    And I have to pull my money out, and that looks bad
    But such cases are rare and one will not make a confident decision for a strict medium rating. For Positive, this album contains lyrics with significant positive values such as Keep Holding On:
    You’re not alone
    Together we stand
    I’ll be by your side, you know I’ll take your hand
    We suppose the density of such obvious expressions is low, but many of the songs are expressing a sense of spirit in pursuit of happiness and love. That makes this album has a medium rating in Positive but hard for the model to capture.
  • Album: A Hard Day’s Night (1964) by The Beatles: The model gives a critical wrong prediction on Positive: predict as low while the ground truth is high. A majority of songs in this album express a deep and genuine emotion about love, such as:
    If you need somebody to love, just look into my eyes
    I’ll be there to make you feel right
    If you’re feeling sorry and sad, I’d really sympathize
    Don’t you be sad, just call me tonight
    The reason that the model made such a mistake could be the lyrics in this album looks plain but the overall expression is obviously positive and constructive.
  • CD single: Labels or Love (2008) by Fergie: The model gives Sex medium, Substance consumption medium, and Consumerism low, while the correct answer are Sex low, Substance consumption low, and Consumerism high. For Sex medium and Substance consumption, the false positive may come from the lexical signals such as sexy, kiss and bag. For Consumerism, there are not only expressions to the encouragement of buying goods:
    Let’s stop chasing those boys and shop some more
    but also many explicit mentions of luxury brands:
    Gucci, Fendi, Prada purses, purchasing them finer things
    Men, they come a dime a dozen
    Just give me them diamond rings
    I’m into a lot of bling, Cadillac, Chanel and Coach
    We suppose the model didn’t take a direct narrative of shopping as a strong signal of Consumerism. Meanwhile, the names of the luxury brand are rare words in the corpus and lack supervision signals. So the model fell hard on this prediction.
6 Ethical considerations and limitations
This work introduces a novel task: assessing positive and risky messages for music products. It also proposes a state-of-the-art method to automatically accomplish the assessment. We acknowledge the potential limitations and ethical considerations by highlighting the following points for future explorations on similar topics:
Reliability: We recognize the potential issues regarding the reliability of such a content rating system. Possible inaccuracies may result in misleading content suggestions, potentially leading vulnerable groups to inadvertently consume inappropriate content, or causing confusion in the production processes for musicians. This work represents our initial exploration, and we strongly advise against implementing such a system in real-world services until the technical and operational elements can be held accountable. We insist that such a system should be regarded as an assistant to, rather than a replacement for, the content rating and assessment work done by media experts and customers.
Social context concerns: The social context in which these labels are acquired is not always known, and there can be a lack of context in terms of why language is used and judged. For instance, many rap songs discuss the harms associated with isolation and substance abuse, yet such information might be misclassified due to rating provider bias or system bias. This may increase the likelihood of systematic bias or unintentionally promote racism. Future research should explore distinguishing between racist and reclaimed uses of slurs.
Data source concerns: Our rating data comes from CSM, a non-neutral organization. The regulations they use to recruit human experts to rate media products are unknown, and different experts rate different products, which may lead to inconsistencies. Transparency and accountability are not guaranteed, and subjectivity remains in the rating results. Future research and implementations should not rashly take ratings from data sources such as CSM as golden standard without careful assessment.
Ambiguity in aspects: The rating aspects used in this work can be ambiguous, as they are loosely defined by single words. For example, what is considered violence may extend beyond overtly violent behaviors. We also recognize that the positive aspect defined by the CSM, indicating the overall takeaway, lacks fine-grained elaborations compared to risky aspects. This could result in ambiguity and fail to provide further insights for users.
Implications and acceptable use: An ethical concern of this study is the potential for media censorship. We acknowledge that efficient machine learning-based algorithms like the proposed method could be used as censorship tools. Malicious users could misuse the proposed method for illegitimate censorship, potentially harming freedom of speech. We call for the development and use of AI algorithms with special attention to who should use the system, how it should be used, and what safeguards should be in place to prevent misuse.
Modality coverage: A technical limitation of this work is that it does not take other modalities of music products, such as melody, rhythm, and vocal performance, into account for predictions. We recognize the significance of these signals in conveying a song’s message. Additionally, we focused only on songs in English in this study, which could underrepresent lyrics written in other languages and from different cultural backgrounds. Future studies should include songs from various languages and cultures.
7 Conclusion
In this paper, we propose a novel task for the NLP community: predicting the strength of aspects in music where we cover objectionable aspects but also ones related to positive content. We leverage lyrics of music products to evaluate various dimensions of messages that they convey to the users. The proposed research problem and approach encourage and motivate further exploration of the music content assessment topic. The presented multi-task ordinality-enforcement model shows significant effectiveness for this task and inspires future research on practical problems with ordinal properties.
Appendix A Implementation details
We list the implementation details of the proposed baseline methods. Sparse and dense document representation: Here we apply TF-IDF and Bag-of-Word-Vectors (GloVe Pennington et al. (2014) Average) as text vectorization approaches and linear models as baselines. TF-IDF method is also used in the previous work that classifies explicitness in lyrics Fell et al. (2020).
Word-level semantic representation: Word vectors such as Word2Vec and Glove are effective semantic representations of NLP tasks. We apply TextCNN Kim (2014) with Glove embedding as a benchmark. The Glove embedding vectors used in the experiments are trained on Wikipedia 2014 and Gigaword 5, with 300 dimensions. The TextCNN has kernel sizes of 3, 4, and 5 in the convolution modules.
Word-level semantic representation with sequence modeling: With word vectors, we further utilize the TextRCNN Lai et al. (2015) model to capture the sequential signal out of each individual word. TextRCNN and other RNN-based models utilize a bi-directional LSTM structure with hidden sizes of 200.
Word-level sequence modeling with attention mechanism: This model performs the best in predicting the MPAA ratings based on movie scripts and rich metadata Shafaei et al. (2020). For script text processing, they apply LSTM to model the sequential information from the word embeddings and use attention mechanism to aggregate the output of each time step for text representation.
Pretrained language model task fine-tuning: Contextualized representation from pretrained Transformer-based models has shown significant success on various NLP tasks. One popular variant, BERT Devlin et al. (2019), was also used in classifying explicitness in the previous workFell et al. (2020). We fine-tune a pretrained BERT on this task in the multi-class classification setting. The BERT model is adapted from HuggingFace with a maximum input length limitation of 512 tokens.
Sentence-level semantic representation with sequence modeling: This model is the state-of-the-art in a severity rating problem for age-restricted content in movies Zhang et al. (2021). We leverage the strong representation capability from the pretrained languages to obtain semantic representations. Then we apply general-purpose sentence embedding from Sentence-BERT Reimers and Gurevych (2019) to encode each sentence from the lyrics of a music item. Then the semantic representation sequences are further encoded using a recurrent architecture to model sequential information. It also becomes a part of the backbone model in our proposed method.
Emotion-guided Transformer model: We apply a Distilled RoBERTa model that is finetuned on emotion detection tasks Hartmann (2022) to obtain the emotion-guided sentence embeddings.
Multitask and ordinality-enforcement: Our proposed multitask model predicts 5 aspects (Positive Messages, Violence, Substance Consumption (Drinking, drugs, and smoking), Sex, and Consumerism) at one single prediction. The ordinality-enforcement components are applied to each individual aspect prediction.
All experiments are conducted using NVIDIA Tesla P40 and PyTorch 1.6.0/PyTorch Lightning 1.0.2. The optimizer is Adam optimizer with 0.001 as the learning rate. Each training epoch of the proposed method takes less than 30 seconds under a batch size of 40.
A.1 Ranking-classification loss behavior
The training behavior of the ranking-classification joint loss in the ordinality-enforcement method is shown in Figure 6. Both cross-entropy losses are averaged on each loss instance in one batch. The ranking loss is often higher during the training process. We suppose ranking is more challenging because the ranking pairs are randomly constructed for every new training step.
Figure 6: Training loss behavior of the ranking-classification method. The classification and ranking losses within one aspect are grouped by the same tone of the color palette: classification losses are in solid lines while ranking losses are in dash lines.